### v1

@@ 2.2 Confined OMP algorithm
  Input: $\boldsymbol{y}\in\mathbb{R}^{m}$, $\boldsymbol{A}\in\{0,1\}^{m\times n}$, $\epsilon$, and $K$.
@@
  Preprocessing: =$\mathcal{E}=\{i:|y_{i}|\leq\epsilon\}$ and $\Gamma=[n]\backslash\bigcup_{i\in\mathcal{E}}\Gamma_{i}^{c}$;
@@
     $t^{(k)}=\underset{i\in\Gamma}{\operatorname*{arg\,max}}|\boldsymbol{A}_{i}^{T}\boldsymbol{r}^{(k-1)}|$(Identification)
@@
We discuss the complexity of OMP algorithm and confined OMP algorithm as follows. It is known that the matrix-vector product can be divided into two steps: multiplication and addition. For a binary-valued matrix, the matrix-vector product only involves addition, no multiplication is required. If non-zero elements of $\boldsymbol{x}$ are real numbers, the addition requires at most $n(d-1)$ float point operations (flops). Furthermore, the selection of the maximum inner-product value in the identification requires $n-1$ flops. Thus, the identification of OMP algorithm requires at most $Kn(d-1)+K(n-1)$ flops.
Thanks to $|\Gamma|\leq n$, the confined OMP algorithm requires fewer inner products in the identification step compared to the conventional OMP algorithm. Specifically, no inner product is required if $|\Gamma|=K$. Otherwise, the identification requires at most $K|\Gamma|(d-1)+K(|\Gamma|-1)$ flops, where $K|\Gamma|(d-1)$ flops are required for the matrix-vector product operations and $K(|\Gamma|-1)$ flops for the selections of the maximum inner-product values. Furthermore, the preprocessing step involves additional operations. Specifically, there are extra $n$ flops to obtain $\mathcal{E}$ and the complexity of obtaining $\Gamma$ is negligible. Consider identification and preprocessing together, these steps requires $n$ flops if $|\Gamma|=K$. Otherwise, at most $K|\Gamma|(d-1)+K(|\Gamma|-1)+n$ flops are required. Table 1 summarizes the complexity of OMP algorithm and confined OMP algorithm in the identification step. The improvement of the proposed algorithm over OMP algorithm is around $Kd$ times in terms of identification efficiency if $|\Gamma|=K$, and $\frac{Knd-K}{K|\Gamma|d-K+n}$ times otherwise.
Table 1: Comparison of the complexity of the OMP algorithm and the confined OMP algorithm in the identification step.
@@
3 Analysis of the proposed algorithm
In this section, we first investigate the expectations of the sparsity of $\boldsymbol{y}$ and the size of the confined set $\Gamma$. Then, we study the lower bounds on the exact recovery probability of OMP algorithm and confined OMP algorithm in the case of using sparse random combinatorial matrices. Note that the lower bound on the exact recovery probability of OMP algorithm is valid for any $K$-sparse signal, whereas that of confined OMP algorithm is only valid for the signal defined in Definition 2.7.
3.1 The sparsity of $\boldsymbol{y}$
Studying the sparsity of $\boldsymbol{y}$ helps to derive the subsequent Lemmas and Theorems. For $k\in\{1,2,\cdots,K\}$, let $\nu^{(k)}=n-|\mathcal{E}|$ be the number of “non-zero” elements of $\boldsymbol{y}=\boldsymbol{A}\boldsymbol{x}$, where the signal $\boldsymbol{x}$ is defined in Definition 2.7 and has $k$ non-zero elements. Obviously, the value of $\nu^{(k)}$ ranges from $d$ to $kd$. The following Lemma gives the probability of $\nu^{(k)}$ for $k=1,2,\cdots,K$.
Lemma 3.1.
@@ 3.1 The sparsity of $\boldsymbol{y}$
(4) $\mathbb{P}\{\nu^{(k)}=\upsilon\}=\sum_{z=\max\{\upsilon-d,d\}}^{\min\{\upsilon,(k-1)d\}}\frac{\binom{z}{\upsilon-z}\binom{m-z}{d-\upsilon+z}}{\binom{m}{d}}\mathbb{P}\{\nu^{(k-1)}=z\}$
for $k=2,3,\cdots,K$ and $\upsilon=d,d+1,\cdots,kd$, where $\mathbb{P}\{\nu^{(k-1)}=z\}$ can be calculated by Eq. 4 recursively.
Proof 3.2.
See Appendix B.
@@
Theorem 3.3 (The expectation of $\nu^{(K)}$).
Suppose that in Eq. 1, $\boldsymbol{A}\in\{0,1\}^{m\times n}$ is a random combinatorial matrix with $d$ ones per column and the signal $\boldsymbol{x}$ is defined in Definition 2.7. Then,
(5) $\mathbf{E}[\nu^{(K)}]=\sum_{\upsilon=d}^{Kd}\upsilon\cdot\mathbb{P}\{\nu^{(K)}=\upsilon\},$
where $\mathbb{P}\{\nu^{(K)}=\upsilon\}$ is given in Lemma 3.1.
Fig. 1 shows empirical means and expectations of $\nu^{(K)}$ for different column degree $d$, where $\boldsymbol{A}\in\{0,1\}^{100\times 256}$ and $\boldsymbol{x}$ is a Gaussian sparse signal with exactly $K$ non-zero elements. It can be seen that the empirical results match well with their expectations.
3.2 The size of $\Gamma$
@@
Theorem 3.4 (The expectation of $|\Gamma|$).
Suppose that in Eq. 1, $\boldsymbol{A}\in\{0,1\}^{m\times n}$ is a random combinatorial matrix with $d$ ones per column and the signal $\boldsymbol{x}$ is defined in Definition 2.7. Then,
(6) $\mathbf{E}[|\Gamma|]=K+(n-K)\cdot\sum_{\upsilon=d}^{Kd}\frac{\binom{\upsilon}{d}}{\binom{m}{d}}\mathbb{P}\{\nu^{(K)}=\upsilon\},$
where $\mathbb{P}\{\nu^{(K)}=\upsilon\}$ is given in Lemma 3.1.
Proof 3.5.
See Appendix C.
@@
Figure 2 shows empirical means and expectations of $|\Gamma|$ for different column degree $d$, where $\boldsymbol{A}\in\{0,1\}^{100\times 256}$ and $\boldsymbol{x}$ is a Gaussian sparse signal with exactly $K$ non-zero elements. It can be seen that the empirical results match well with their expectations. We also observe that $\mathbf{E}[|\Gamma|]\approx K$ when $K$ is relatively small. These observations confirm the effectiveness of removing the redundancy of $\boldsymbol{A}$.
3.3 Recovery performance analysis of OMP
@@
In this subsection, we provide a lower bound on the probability that the OMP algorithm exactly recovers any $K$-sparse signal $\boldsymbol{x}$ in $K$ iterations for $\boldsymbol{A}\in\{0,1\}^{m\times n}$.
We first define the coherence between two columns of $\boldsymbol{A}$ as
(7) $\mu_{i,j}=\frac{|\boldsymbol{A}_{i}^{T}\boldsymbol{A}_{j}|}{\lVert\boldsymbol{A}_{i}\rVert_{2}\lVert\boldsymbol{A}_{j}\rVert_{2}},$
where $i,j\in[n]$ and $i\neq j$. Different from the random Gaussian matrices [40], the values of $\mu_{i,j}$ are discrete for the binary-valued matrix $\boldsymbol{A}$. That is, we have $\mu_{i,j}\in\{0,1/d,2/d,\cdots,1\}$. We further use $g(s)$ for $s\in\{0,1/d,2/d,\cdots,1\}$ to denote the distribution function of $\mu_{i,j}$:

### v2

@@ II-B Confined OMP Algorithm
 Input: $\boldsymbol{y}\in\mathbb{R}^{m}$, $\boldsymbol{A}\in\{0,1\}^{m\times n}$, $\epsilon$, $\varepsilon$, and $K$.
@@
[…]
@@
  $t^{(k)}=\underset{i\in\Gamma\backslash\Lambda^{(k-1)}}{\operatorname*{arg\,max}}|\boldsymbol{A}_{i}^{T}\boldsymbol{r}^{(k-1)}|$(Identification)
@@
We discuss the complexity of OMP algorithm and confined OMP algorithm as follows. It is known that the matrix-vector product can be divided into two steps: multiplication and addition. For a binary-valued matrix, the matrix-vector product only involves addition, no multiplication is required. If nonzero elements of $\boldsymbol{x}$ are real numbers, the addition requires at most n(d1)n(d-1) floating-point operations (flops). Furthermore, the selection of the maximum inner-product value in the identification requires n1n-1 flops. Thus, the identification of OMP algorithm requires at most Kn(d1)+K(n1)Kn(d-1)+K(n-1) flops.
Thanks to $|\Gamma|\leq n$, the confined OMP algorithm requires fewer inner products in the identification step compared to the conventional OMP algorithm. Specifically, no inner product is required if $|\Gamma|=K$. Otherwise, the identification requires at most $K|\Gamma|(d-1)+K(|\Gamma|-1)$ flops, where $K|\Gamma|(d-1)$ flops are required for the matrix-vector product operations and $K(|\Gamma|-1)$ flops for the selections of the maximum inner-product values. Furthermore, the preprocessing step involves additional operations. Specifically, there are extra $n$ flops to obtain $\mathcal{E}$ and the complexity of obtaining $\Gamma$ is negligible. Consider identification and preprocessing together, these steps require $n$ flops if $|\Gamma|=K$. Otherwise, at most $K|\Gamma|(d-1)+K(|\Gamma|-1)+n$ flops are required. Table I summarizes the complexity of OMP algorithm and confined OMP algorithm in the identification step. The improvement of the proposed algorithm over the OMP algorithm is around $Kd$ times in terms of identification efficiency if $|\Gamma|=K$, and $\frac{Knd-K}{K|\Gamma|d-K+n}$ times otherwise.
TABLE I: Comparison of the complexity of the OMP algorithm and the confined OMP algorithm in the identification step.
@@
II-C Extension to the Generalized OMP Algorithm
One advantage of the confined set is that it can be easily applied to many other sparse signal recovery algorithms. In this subsection, we use the generalized OMP (gOMP) algorithm as an example for illustration.
The gOMP algorithm [14, 13] is an efficient greedy recovery algorithm that allows selecting multiple column indices maximally correlated with the residual at each iteration, thereby reducing the required number of iterations. Similar to the confined OMP algorithm, the confined gOMP algorithm is a variant of gOMP algorithm that confines the support estimation to a reduced subset $\Gamma$ of the dictionary. Let $N$ be the number of column indices selected by the confined gOMP algorithm in the identification step of each iteration. The details of the confined gOMP algorithm are given in Algorithm 3. The main differences are summarized below.
  • (Preprocessing) If $|\Gamma|\leq\max\{K,N\}$, then the confined gOMP algorithm directly performs the least squares to estimate the sparse signal $\boldsymbol{x}$. Compared to the confined OMP algorithm, the confined gOMP algorithm allows for some extra redundancy in the preprocessing step when $K<N$.
  • (Identification) In the identification step, if $N\leq|\Gamma\backslash\Lambda^{(k-1)}|$, column indices corresponding to the largest $N$ correlation in magnitude are selected to form a set $\mathcal{T}^{(k)}$. Otherwise, the remaining column indices in $\Gamma\backslash\Lambda^{(k-1)}$ are chosen.
  • (Stopping criteria) The iterative process is terminated if $\lVert\boldsymbol{r}^{(k)}\rVert_{2}\leq\varepsilon$, the iteration number reaches maximum $\min\{K,\lfloor\frac{m}{N}\rfloor\}$, or the size of the estimated support set $\Lambda^{(k)}$ is equal to $|\Gamma|$.
Algorithm 3 Confined generalized orthogonal matching pursuit algorithm
 Input: $\boldsymbol{y}\in\mathbb{R}^{m}$, $\boldsymbol{A}\in\{0,1\}^{m\times n}$, $\epsilon$, $\varepsilon$, $N$, and $K$.
[…]
if $|\Gamma|\leq\max\{K,N\}$ then
  return $\hat{\boldsymbol{x}}$.
end if
[…]
  $k=k+1$,
  if $N\leq|\Gamma\backslash\Lambda^{(k-1)}|$ then
   $\mathcal{T}^{(k)}=\underset{\begin{subarray}{c}|\mathcal{T}^{(k)}|=N\\ i\in\Gamma\backslash\Lambda^{(k-1)}\end{subarray}}{\operatorname*{arg\,max}}|\boldsymbol{A}_{i}^{T}\boldsymbol{r}^{(k-1)}|$(Identification)
  else
   $\mathcal{T}^{(k)}=\Gamma\backslash\Lambda^{(k-1)}$(Identification)
  end if
  $\Lambda^{(k)}=\Lambda^{(k-1)}\cup\mathcal{T}^{(k)}$(Augmentation)
  $\hat{\boldsymbol{x}}_{\Lambda^{(k)}}^{(k)}=\underset{\boldsymbol{x}\in\mathbb{R}^{|\Lambda^{(k)}|}}{\operatorname*{arg\,min}}\lVert\boldsymbol{y}-\boldsymbol{A}_{\Lambda^{(k)}}\boldsymbol{x}\rVert_{2}$(Estimation)
end while
return $\hat{\boldsymbol{x}}^{(K)}$.
III Analysis of The Proposed Algorithm
In this section, we first investigate the expectations of the sparsity of $\boldsymbol{y}$ and the size of the confined set $\Gamma$. Then, the recovery performance of the confined OMP algorithm using sparse random combinatorial matrices is analyzed under a noiseless linear system. Finally, we investigate the recovery performance and robustness of the proposed algorithm in a noisy linear system.
III-A The Sparsity of $\boldsymbol{y}$
Studying the sparsity of $\boldsymbol{y}$ helps to derive the subsequent Lemmas and Theorems. Let $\epsilon\to 0$ and $\mathcal{E}=\{i:|y_{i}|\leq\epsilon\}$. Then, for $k\in\{1,2,\cdots,K\}$, let $\nu^{(k)}=m-|\mathcal{E}|$ be the number of “nonzero” elements of $\boldsymbol{y}=\boldsymbol{A}\boldsymbol{x}$, where $\boldsymbol{x}$ is the confined signal and has $k$ nonzero elements. Obviously, the value of $\nu^{(k)}$ ranges from $d$ to $\min\{kd,m\}$. The following Lemma gives the probability of $\nu^{(k)}$ for $k=1,2,\cdots,K$.
Lemma 1
@@ III-A The Sparsity of $\boldsymbol{y}$
$\mathbb{P}\{\nu^{(k)}=\upsilon\}=\sum_{z=\max\{\upsilon-d,d\}}^{\min\{\upsilon,(k-1)d\}}\frac{\binom{z}{\upsilon-z}\binom{m-z}{d-\upsilon+z}}{\binom{m}{d}}\mathbb{P}\{\nu^{(k-1)}=z\}$ (4)
for $k=2,3,\cdots,K$ and $\upsilon=d,d+1,\cdots,\min\{kd,m\}$, where $\mathbb{P}\{\nu^{(k-1)}=z\}$ can be calculated by (4) recursively.
Proof:
See Appendix B. ∎
@@
Lemma 2 (The expectation of $\nu^{(K)}$)
Suppose that in (1), $\boldsymbol{A}\in\{0,1\}^{m\times n}$ is a random combinatorial matrix with $d$ ones per column and the signal $\boldsymbol{x}$ is a confined signal. Then,
$\mathbf{E}_{\upsilon}[\nu^{(K)}]=\sum_{\upsilon=d}^{\min\{Kd,m\}}\upsilon\cdot\mathbb{P}\{\nu^{(K)}=\upsilon\},$ (5)
where $\mathbb{P}\{\nu^{(K)}=\upsilon\}$ is given in (4). Also, it can be easily calculated by
$\mathbf{E}_{\upsilon}[\nu^{(K)}]=m\left(1-\left(1-\frac{d}{m}\right)^{K}\right).$ (6)
Fig. 1 shows empirical means and expectations of $\nu^{(K)}$ for different column degree $d$, where $\boldsymbol{A}\in\{0,1\}^{100\times 256}$ and $\boldsymbol{x}$ is a Gaussian sparse signal with exactly $K$ nonzero elements. It can be seen that the empirical results match well with their expectations.
@@
III-B The Size of $\Gamma$
@@
Theorem 2 (The expectation of $|\Gamma|$)
Suppose that in (1), $\boldsymbol{A}\in\{0,1\}^{m\times n}$ is a random combinatorial matrix with $d$ ones per column and the signal $\boldsymbol{x}$ is a confined signal. Then,
$\mathbf{E}[|\Gamma|]=K+(n-K)\cdot\mathbf{E}_{\upsilon}\left[\frac{\binom{\nu^{(K)}}{d}}{\binom{m}{d}}\right$], (7)
where $\mathbf{E}_{\upsilon}[\cdot]$ is given in (5).
Proof:
See Appendix C. ∎
@@
Fig. 2 shows empirical means and expectations of $|\Gamma|$ for different column degree $d$, where $\boldsymbol{A}\in\{0,1\}^{100\times 256}$ and $\boldsymbol{x}$ is a Gaussian sparse signal with exactly $K$ nonzero elements. It can be seen that the empirical results match well with their expectations. We also observe that $\mathbf{E}[|\Gamma|]\approx K$ when $K$ is relatively small. These observations confirm the effectiveness of removing the redundancy of $\boldsymbol{A}$.
III-C Analysis of Confined OMP in the Noiseless Linear System
We define $spark(\boldsymbol{A})$ as the spark of $\boldsymbol{A}$, where $spark(\boldsymbol{A}):=\min\{\lVert\boldsymbol{x}\rVert_{0}:\boldsymbol{A}\boldsymbol{x}=\boldsymbol{0},\boldsymbol{x}\neq\boldsymbol{0}\}$. The condition $spark(\boldsymbol{A})>K$ is essential to guarantee that the least squares problem in the proposed algorithm has a unique solution. The spark of $\boldsymbol{A}$ depends heavily on the column degree $d$. For example, when the column degree is set to $d=1$, we have $spark(\boldsymbol{A})=2$. However, the computation of $spark(\boldsymbol{A})$ is NP-hard in general. Fortunately, the following Theorem and Corollary show that when $m$ is sufficiently large, $\boldsymbol{A}$ meets $spark(\boldsymbol{A})>K$ with high probability.
Theorem 3 (Theorem 1.2 in [38])
Fix $\gamma>0$, and let $d=d(m)$ be any function of $m$ satisfying $\min\{d,m-d\}\geq(1+\gamma)\ln m$. Then, for an $m\times m$ random combinatorial matrix $\boldsymbol{Q}$ with independent rows, where each row is chosen uniformly among the vectors with $d$ ones, we have
Corollary 3
Fix $\gamma>0$, and let $d=d(m)$ be any function of $m$ satisfying $(1+\gamma)\ln m\leq d\leq m/2$. Then, for an $m\times n$ random combinatorial matrix $\boldsymbol{A}$ with $d$ ones per column and a sufficiently large $m$, we have
$\mathbb{P}\left\{spark(\boldsymbol{A})>K\right\}=1-o(1).$ (8)
Proof:
According to Theorem 3, we know that the $m\times m$ random combinatorial matrix $\boldsymbol{Q}$ is nonsingular with probability 1o(1)1-o(1) for a sufficiently large $m$. In other words, $m$ columns of $\boldsymbol{Q}^{T}$ are linearly independent with probability 1o(1)1-o(1) for a sufficiently large $m$. For the $m\times n$ random combinatorial matrix $\boldsymbol{A}$ with a sufficiently large $m$, any $K$ columns of $\boldsymbol{A}$ are obviously linearly independent with probability 1o(1)1-o(1), resulting in $spark(\boldsymbol{A})>K$ with probability 1o(1)1-o(1). Thus, we get (8). ∎
In general, the sparsity level $K$ is much smaller than $m$. Thus, as $m$ increases, the probability $\mathbb{P}\{spark(\boldsymbol{A})>K\}$ converges to 1 more rapidly than the probability $\mathbb{P}\{\boldsymbol{Q}~\mathrm{is}~\mathrm{nonsingular}\}$ shown in Theorem 3. In summary, when $m$ is sufficiently large and the column degree $d$ satisfies $(1+\gamma)\ln m\leq d\leq m/2$ for a constant $\gamma>0$, the spark of $\boldsymbol{A}$ is greater than $K$ with probability 1o(1)1-o(1). In the following theoretical analysis, we will take this condition as a premise.
@@ III-C Analysis of Confined OMP in the Noiseless Linear System
Define the event $\mathbb{S}_{\mathrm{comp}}=\{$confined OMP exactly recovers the $K$-sparse confined signal $\boldsymbol{x}\}$. In the following, we will present a lower bound on the probability that the confined OMP algorithm exactly recovers the confined signal $\boldsymbol{x}$ for $\boldsymbol{A}\in\{0,1\}^{m\times n}$.
Theorem 4 (Recovery probability of confined OMP)
Suppose that in (1), $\boldsymbol{A}$ is an $m\times n$ random combinatorial matrix with $d$ ones per column, and $\boldsymbol{x}$ is a confined signal with exactly $K$ nonzero entries. If $\boldsymbol{A}$ satisfies $spark(\boldsymbol{A})>K$, then it holds that